OpenStudy (anonymous):

The ratio of two numbers is 1 to 5, and the sum of the numbers is 36. What is the smaller number? (Points : 4) 6 7 8 9

OpenStudy (anonymous):

@uri @StudyGurl14

OpenStudy (anonymous):

@DarkBlueChocobo

OpenStudy (freckles):

let x and y be the two numbers mentioned

OpenStudy (freckles):

you are given the following \[\frac{x}{y}=\frac{1}{5} \\ x+y=36\]

OpenStudy (anonymous):

ya

OpenStudy (uri):

1:5 Total :36 Ratio of smaller number? 1+5=6 1/6*36

OpenStudy (uri):